Solve:(p − q)z = z^2 + (x + y)^2
1
Expert's answer
2020-10-04T18:41:39-0400

Given equation is "(p-q)z = z^2 + (x+y)^2"

Comparing above equation with "Pp + Qq = R"

Then

"\frac{dx}{z}=\frac{dy}{-z}=\frac{dz}{z^2 + (x+y)^2}"


Taking first two,

"\frac{dx}{z} = - \frac{dy}{z}"

Integrating both sides,

"x = - y + C"

"x+y = C" (1)


Then taking first and last,

"\frac{dx}{z} = \frac{dz}{z^2+(x+y)^2}"


"\frac{dx}{z} = \frac{dz}{z^2+C^2}"

Integrating both sides


"\int {dx}=\int \frac{zdz}{z^2+C^2}"


"x = \frac{1}{2}log|z^2 + C^2| + log|b|"

then "b = (z^2 + (x+y)^2)e^{-2x}"


Hence,

solution of the equation will be

"f(x+y) = (z^2 + (x+y)^2)e^{-2x}"